OpenStudy (anonymous):

Write an equation of the line that passes through (8, 3) and (4, 3).

OpenStudy (lgbasallote):

use the formula \[y - y_1 = \frac{y_2 - y_1}{x_2 - x_1} ( x- x_1)\] do you know how to use it?

OpenStudy (anonymous):

wouldn't it just be y2-y1/x2-x1 w/o those other things.

OpenStudy (lgbasallote):

that's the slope

OpenStudy (anonymous):

o, then no idk how to use it.

OpenStudy (lgbasallote):

but you know how to use y2-y1/x2-x1?

OpenStudy (anonymous):

ya.

OpenStudy (lgbasallote):

okay good.. solve for that. then tell me what you get

OpenStudy (anonymous):

I got 0/4

OpenStudy (lgbasallote):

so simplify it...what do you get?

OpenStudy (anonymous):

0

OpenStudy (lgbasallote):

good so like i said that's the slope

OpenStudy (lgbasallote):

now do you know the point-slope formula \[y - y_1 = m(x-x_1)\]

OpenStudy (anonymous):

no, I dont understand it.

OpenStudy (lgbasallote):

okay..look back in what you did for \[\frac{y_2 - y_1}{x_2 - x_1}\] what did you use as y1?

OpenStudy (anonymous):

3

OpenStudy (lgbasallote):

and what did you use as x1?

OpenStudy (anonymous):

8

OpenStudy (lgbasallote):

so just put that into the equation i gave. like i said m is 0 (you solved this yourself) \[y - y_1 = m(x-x_1)\] \[y - 3 = 0(x-8)\]

OpenStudy (lgbasallote):

simplify this now

OpenStudy (anonymous):

Got 3.